OpenStudy (anonymous):

Use each distributive property to simplify each math problem (a+b)^6+3(2a+b)

OpenStudy (anonymous):

(a + b)^6? Have you learned the binomial theorem yet? Otherwise this problem is tedious.

OpenStudy (anonymous):

I am in 6th grade and don't know this problem. Those words are pretty big.

OpenStudy (anonymous):

Hmm... Did you copy the problem correctly? (a + b) raised to the power 6? That's pretty big. Maybe there's a typo in the book.

OpenStudy (anonymous):

It's a worksheet she said she wanted us to do it or get help doing it and i have no clue on how to do it.

OpenStudy (anonymous):

Oh wow. OK (a + b)^6 means (a + b)(a + b)(a + b)(a + b)(a + b)(a + b). So take it one at a time. (a + b)(a + b) = ? Can you FOIL it out?

OpenStudy (anonymous):

8a+8b=?

OpenStudy (anonymous):

Not quite. Let's try this. (x + 1)(x + 1) = x(x + 1) + 1(x + 1). So (x + 1)(x + 1) = x^2 + x + x + 1 = x^2 + 2x + 1. So (a + b)(a + b) = a(a + b) + b(a + b) = a^2 + ab + ba + b^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

(x + 1)(x + 1) = x(x + 1) + 1(x + 1), that's distributing the multiplication.

OpenStudy (anonymous):

I'm not sure what your teacher wants you to do. Doing the multiplication out directly seems like busywork.